### v1

Abstract
We extend a previous result [Sutter et al, J. Phys. A: Math. Theor. 57, 475202 (2024)] to give an explicit form of the set of $v$-representable densities on the one-dimensional torus with any fixed number of particles in contact with a heat bath at finite temperature. The particle interaction has to satisfy some mild assumptions but is kept entirely general otherwise. For densities, we consider the Sobolev space $H^{1}$ and exploit the convexity of the functionals. This leads to a broader set of potentials than the usual $L^{p}$ spaces and encompasses distributions. By including temperature and thus considering all excited states in the Gibbs ensemble, Gâteaux differentiability of the thermal universal functional is guaranteed. This yields $v$-representability and it is demonstrated that the given set of $v$-representable densities is even maximal.
I Introduction
Density-functional theory (DFT) is based on the one-to-one relation between densities and potentials. The famous work of Hohenberg and Kohn [1] introduced an energy functional in terms of the density which as such is only well-defined for densities coming from a ground state, a property called ‘pure-state $v$-representability’. The lack of knowledge about this property can be circumvented by using the constrained-search functional as introduced by Percus [2] and Levy [3]. This functional is defined for all densities coming from a $N$-particle wave function ($N$-representability). However, for the method of Kohn and Sham [4] it is still crucial to solve the $v$-representability problem. The idea of this method is to reproduce the ground-state density of an interacting Hamiltonian by a non-interacting Hamiltonian with another external potential. Thus, the density needs to be $v$-representable for an interacting system but at the same time also for a non-interacting one. A possible extension is to allow for mixed states. Levy [5] and Lieb [6] showed that there are ensemble ground-state density which can not be realized by a pure ground-state. Hence the set of $v$-representable densities is extended by allowing for ensemble states and not only pure states. In [7], it is shown that each strictly positive density in $H^{1}(\mathbb{T})$ on the one-dimensional torus can be realized by a distributional potential in $H^{-1}(\mathbb{T})$. Shortly after, Carvalho Corso [8, 9] showed that for an odd number of particles on the one-dimensional torus each ground-state density of a Hamiltonian with a potential in $H^{-1}(\mathbb{T})$ is non-zero. This result completes the full characterization of $v$-representable densities on the one-dimensional torus. Apart from that it was shown that under the same condition for the number of particles, the external potential is uniquely determined by the ground-state density and that the constrained-search functional is differentiable at strictly positive densities. Additionally, Carvalho Corso [10] presented these results for non-interacting systems but without any requirement on the number of particles.
Already in 1965 Mermin [11] extended the work of Hohenberg and Kohn [1] to non-zero temperatures. However the same question about $v$-representability appears. Therefore we are interested in extending our recent work [7] from zero temperature to elevated temperatures. First results about this problem were presented by Chayes et al. [12] for infinite lattice systems. Eschrig [13] formally argued that for the grand canonical ensemble on a $3$-dimensional torus the thermal universal functional is differentiable at $v$-representable densities and that these densities are strictly positive. However, he does not give an explicit expression for the set of $v$-representable densities. In Eschrig’s work, the underlying topology is the usually considered Lebesgue space $L^{3}$, yet the domain of the universal functional (the set of physical densities) is not open in $L^{3}$. In particular, arbitrary close to any physical density there is n density which takes negative values. Hence the universal functional is discontinuous at each physical density. This problem is already extensively studied by Lammert [14] for the zero temperature case. While Lammert is able to show differentiability at certain densities, he does not characterize the set of $v$-representable densities, a problem finally solved on the one-dimensional torus in Sutter et al. [7]. In our previous and the present work, the problems with openness and continuity are circumvented by considering the $H^{1}$ topology instead of the $L^{3}$ topology. Then, the set of $v$-representable densities is open and the universal functional is continuous on that set. Hence, we can investigate Gâteaux differentiability. Another important difference to Eschrig’s work is that we allow for a broader class of potentials and we show that this is even necessary for $v$-representability.
In this work we consider the same class of systems as in Sutter et al. [7] but at elevated temperatures. This means that there are a fixed number of particles on the one-dimensional torus and that we allow for a whole class of interactions including $W=0$. Elevated temperature means that besides the Hamiltonian an additional entropic term has to be considered in the variational principle. In particular, this means that instead of the usual internal energy functional we work with the Helmholtz functional. The minimizer of the Helmholtz functional is explicitly known in terms of the Hamiltonian and it is called Gibbs state. As in [15; 16] the entropy term acts as a regularizer and we are not only able to show that each strictly positive density is $v$-representable but we also get (Gâteaux) differentiability of the universal functional for all interactions in the considered class. Additionally, we show that each density coming from a Gibbs state is strictly positive as Carvalho Corso [8, 9] showed for zero temperature and a odd number of particles. Hence we get the full characterization of $v$-representable densities. The reason for these additional results compared to the zero temperature case with even number of particles, is that the minimizer of Helmholtz functional occupies not only the ground state of the Hamiltonian but also all excited states. Degenerate states get equal weights which yields a unique minimizer of the Helmholtz functional and strict convexity of the Helmholtz free energy functional (in terms of potentials). We formulate the main result of this work.
Theorem 1.
Define the following set of densities on the one-dimensional torus $\mathbb{T}$ and the space of (distributional) one-body potentials,
$\mathscr{X}_{>0}$ $=\{\rho\in L^{2}(\mathbb{T})\mid\nabla\rho\in L^{2}(\mathbb{T}),\smallint\rho=N,\forall x\in\mathbb{T}:\rho(x)>0\}$ (1a)
$\mathscr{X}^{*}$ $=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (1b)
Then for every $\rho\in\mathscr{X}_{>0}$ there is an equivalence class $v]\in\mathscr{X}^{*}$ with corresponding potential $V=\sum_{j=1}^{N}v(x_{j})$, such that $\rho$ is the density of the Gibbs state (see Proposition 3) corresponding to the self-adjoint Hamiltonian $H_{v}=H_{0}+V=-\frac{1}{2}\Delta+W+V$, with $W$ any (interaction) operator that fulfills the KLMN condition (see Eq. (3)). On the other hand, each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. Moreover, the thermal universal functional $F^{\beta}_{\mathrm{DM}}(\rho)$ (defined in Eq. (15b)) is differentiable if and only if $\rho\in\mathscr{X}_{>0}$. This implies that every $\rho\in\mathscr{X}_{>0}$ is even uniquely $v$-representable by a $v]\in\mathscr{X}^{*}$.
The ‘unique $v$-representability’ is equivalent to the famous Hohenberg–Kohn theorem [1]. To prove the stated theorem above, we start by introducing the basic setting for a $N$-particle system on a one-dimensional torus and the relevant spaces for the densities and the potentials (Section II). Afterwards, we introduce the Helmholtz functional. We present known result about this functional and about the Gibbs state together with the introduction of the thermal universal functional (Section III). Then we show the relation between $v$-representability and subgradients (Section IV) followed by an application of the preceding results to conclude $v$-representability (Section V). We then prove the converse statement, that each Gibbs state has indeed a density in the $v$-representable set, to get a full characterization of the set of $v$-representable densities (Section VI). Lastly, we show that the universal functional is Gâteaux differentiable (Section VII). We conclude with the proof of Theorem 1 and an outlook (Section VIII).
II One-dimensional torus setting
We consider a system at an elevated temperature $T$ consisting of $N$ spin $\frac{1}{2}$-particles. The particles are on the one-dimensional torus $\mathbb{T}$, meaning that they are confined to the interval $[0,1]$ where the points 0 and $1$ are identified with each other. The underlying Hilbert space is then
$\mathcal{H}=\bigl{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}.$ (2)
As noted, the considered Hamiltonians are of the form $H_{v}=H_{0}+V$ where $H_{0}$ consists of the kinetic energy operator and an interaction term and where $V$ is the potential. We can generally write $H_{v}$ instead of $H_{[v]}$, even though the considered potentials come from equivalence classes, since adding a constant to the potential just shifts the eigenvalues while not affecting the eigenstates. The interaction is kept general but positive and it needs to fulfill the KLMN condition (Reed and Simon [17], Th. X.17) with respect to the kinetic energy operator, i.e. for each interaction $W$ there are $0\leq a<1$ and $b\geq 0$ such that the following holds for each element in the form domain $Q(-\frac{1}{2}\Delta)$
$|\langle\Psi,W\Psi\rangle|\leq a\langle\Psi,-\tfrac{1}{2}\Delta\Psi\rangle+b\langle\Psi,\Psi\rangle.$ (3)
This condition guarantees that the Hamiltonian $H_{0}=-\tfrac{1}{2}\Delta+W$ remains self-adjoint with the same form domain $Q(H_{0})=Q(-\frac{1}{2}\Delta)$ if we add an interaction to the kinetic energy operator (which is self-adjoint). In Sutter et al. [7, Appendix] this condition was shown for interactions in $L^{1}([-1,1])$ and for distributional interactions of the form $\nabla g$ with $g\in L^{2}([-1,1])$. Here and in the following we use the notion of quadratic forms [18, Sec. VIII.6] instead of linear operators. For the kinetic energy operator the form domain is given by $Q(-\frac{1}{2}\Delta)=\bigl{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\bigr{)}^{\wedge N}$ with the Sobolev space $H^{1}(\mathbb{T})=\{\Psi\in L^{2}(\mathbb{T})\mid\nabla\Psi\in L^{2}(\mathbb{T})\}$ (like Eq. 1a this means a weak derivative). We get by integration by parts
$\frac{1}{2}\langle\Psi,-\Delta\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\Psi,-\Delta_{j}\Psi\rangle=\frac{1}{2}\sum_{j=1}^{N}\langle\nabla_{j}\Psi,\nabla_{j}\Psi\rangle.$ (4)
Together with the normalization constraint this gives the set of wave functions
$\mathscr{W}=\{\Psi\in(H^{1}(\mathbb{T})\otimes\mathbb{C}^{2})^{\wedge N}\mid\|\Psi\|_{\mathcal{H}}=1\}.$ (5)
The one-particle density $\rho_{\Psi}$ is obtained by integrating out all but one of the spatial coordinates of a wave function squared as well as summing over all spin components, i.e.
$\rho_{\Psi}(x)=N\sum_{\sigma_{1},\ldots,\sigma_{N}}\int\lvert\Psi(x\sigma_{1},x_{2}\sigma_{2},\dotsc,x_{N}\sigma_{N})\rvert^{2}\mathrm{d}x_{2}\dotsb\mathrm{d}x_{N}.$ (6)
In the following theorem we collect results from [6] and [7].
Theorem 2.
We define the following set of densities
$\mathscr{I}=\{\rho\in L^{1}(\mathbb{T})\mid\nabla\sqrt{\rho}\in L^{2}(\mathbb{T}),\rho\geq 0,\smallint\rho=N\}.$ (7)
Then, the following holds.
  1. (i)
    Let $\Psi\in\mathscr{W}$. Then $\rho_{\Psi}\in\mathscr{I}$ and it holds
    $\int\left|\nabla\sqrt{\rho_{\Psi}(x)}\right|^{2}\mathrm{d}x\leq 2T(\Psi).$ (8)
  2. (ii)
    There exist two constants $C_{1},C_{2}>0$ such that for all $\rho\in\mathscr{I}$ there is a $\Psi\in\mathscr{W}$ that has $\rho_{\Psi}=\rho$ and the following estimate holds
    $T(\Psi)\leq C_{1}+C_{2}\int\left|\nabla\sqrt{\rho(x)}\right|^{2}\mathrm{d}x.$ (9)
  3. (iii)
    $\mathscr{X}_{>0}\subset\mathscr{I}\subset\mathscr{X}\subset H^{1}(\mathbb{T})\hookrightarrow C^{0}(\mathbb{T})\hookrightarrow L^{\infty}(\mathbb{T})\hookrightarrow L^{3}(\mathbb{T})\hookrightarrow L^{2}(\mathbb{T})\hookrightarrow L^{1}(\mathbb{T})$, where $\mathscr{X}=\{\rho\in H^{1}(\mathbb{T})\mid\smallint\rho=N\}$ and where $X\hookrightarrow Y$ means that $X$ is continuously embedded in $Y$.
Lieb [6] showed that the set $\mathscr{I}$ is convex. This implies that (finite) convex combinations of elements in $\mathscr{I}$ lie again in $\mathscr{I}$. Motivated by the observations for the zero temperature case [7] we identify the space of external potential by $\mathscr{X}^{*}$, the topological dual of $\mathscr{X}$. This space contains equivalence classes, where the elements are from the space $H^{-1}(\mathbb{T})$ and where two elements are in the same class if they differ only by a constant, i.e., like in Eq. (1b),
$\mathscr{X}^{*}=\{[v]=\{v+c\mid c\in\mathbb{R}\}\mid v=f+\nabla g\;\text{with}\;f,g\in L^{2}(\mathbb{T})\}.$ (10)
The dual pairing of an element $v=f+\nabla g\in H^{-1}(\mathbb{T})$ with an element $\varphi\in H^{1}(\mathbb{T})$ is given by $v(\varphi)=\langle v,\varphi\rangle=\langle f,\varphi\rangle-\langle g,\nabla\varphi\rangle$. Using the dual space of $\mathscr{X}$ for the external potentials guarantees that the potential energy is finite. That Hamiltonians of the form $H=H_{0}+V$ with $V$ defined by a $v\in\mathscr{X}^{*}$ are connected to a self-adjoint operator is shown in [7] by using the KLNM theorem (Reed and Simon [17], Th. X.17).
III Helmholtz free energy and Gibbs state
Considering a system at fixed temperature $T>0$ (through coupling to a heat bath) means that we cannot rely on a variational principle for the internal energy $E_{v}(\Gamma)=\operatorname{Tr}\{\Gamma H_{v}\}$ any more [19]. Instead, we have to consider the Helmholtz free energy
$\Omega_{v}^{\beta}(\Gamma)=E_{v}(\Gamma)-TS(\Gamma)=\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}},$ (11)
where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The new variational principle then yields the Helmholtz functional
$\Omega^{\beta}(v)=\inf_{\Gamma}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma}\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}}$ (12)
and can been equivalentely seen as the Legendre transform of $E_{v}$ (in terms of entropy) to $\Omega^{\beta}(v)$ (in terms of temperature). Here the infimum is over all density matrices, i.e. over all (possibly infinite) convex sums $\Gamma=\sum_{k}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}$, $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$, with $\Psi_{k}\in\mathscr{W}$ all orthogonal. Making the Helmholtz functional stationary with respect to small perturbation $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
Proposition 3.
The unique minimizer of the Helmholtz functional is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function. The state $\Gamma_{v}$ is called the Gibbs state and it does only exist if $Z(v)$ is finite. Moreover, for each $v\in H^{-1}(\mathbb{T})$ the Gibbs state exists and the Helmholtz free energy is given by
$\Omega^{\beta}(v)=-\beta^{-1}\log\bigl{(}Z(v)\bigr{)}.$ (13)
The proof of Proposition 3 can be found in Appendix A. The reason for $Z(v)$ to be finite is that the particles are trapped in a finite volume. The defining parameters of canonical (quantum) thermodynamics are $(N,V,T)$ and all of these parameters need to be finite. For a counterexample [16, Sec. 2.3] consider the hydrogen atom in $\mathbb{R}^{3}$. The energies are $1/(2n^{2})$ with a degeneracy of $n^{2}$ and with $n\in\mathbb{N}$. Then the partition function is
$Z(v)=\sum\limits_{n=1}^{\infty}n^{2}\mathrm{e}^{\beta/(2n^{2})}\geq\sum\limits_{n=1}^{\infty}n^{2}=\infty.$ (14)
Proposition 3 implies that the density of the Gibbs state is an infinite convex combination of elements in $\mathscr{I}$. Hence we cannot directly conclude that its density is contained in $\mathscr{I}$, but we will show this statement with the following proposition.
Proposition 4.
The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}$.
The proposition is proven in Appendix B. It tells us that $v$-representable densities at elevated temperatures are contained in $\mathscr{I}$. Hence we can split the minimization of the Helmholtz functional $\Omega^{\beta}_{v}$ as in the zero temperature case into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density,
$\Omega^{\beta}(v)=\inf_{\rho\in\mathscr{I}}\{F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle\}\quad\text{where}$ (15a)
$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\Gamma\mapsto\rho}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (15b)
The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$. Note that because of the existence of the Gibbs state, the infimum in the variation over the densities is attained whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
As for zero temperature ([6; 7]), we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions. Due to the concavity of the entropy [16, Th. 12] and linearity of $\Gamma\to\rho$ the following proposition follows.
Proposition 5.
The following two statements hold.
  1. (i)
    The Helmholtz functional $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex.
  2. (ii)
    The universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
IV Relation between $v$-representability and subdifferentials
The relation between $v$-representability and non-empty subdifferential was presented in the literature [20; 21; 22]. This relation, together with some additional results from convex analysis, are important ingredients to show our main result Theorem 1. Therefore, we will repeat the relation and the needed results in this section.
For any proper ($f\neq\infty$) convex function $f:\mathscr{X}\to(-\infty,\infty]$ its subdifferential at a point $\rho\in\mathscr{X}$ is defined as the set
$\partial f(\rho)=\{u\in\mathscr{X}^{*}\mid f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle\;\forall\rho^{\prime}\in\mathscr{X}\}.$ (16)
Note, that if the inequality $f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle$ cannot be satisfied for any $u\in\mathscr{X}^{*}$ and all $\rho^{\prime}\in\mathscr{X}$ then the subdifferential is empty. An element of $\partial f(\rho)$ is called a subgradient. Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [23, Prop. 2.40]. In this case, this element is equal to the derivative. If the general convex function is replaced by the universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)+\langle v,\rho\rangle\leq F_{\mathrm{DM}}^{\beta}(\rho^{\prime})+\langle v,\rho^{\prime}\rangle.$ (17)
In particular, $\rho$ is a minimizer for the variation in Eq. (15a). Note that this is only a necessary condition for $v$-representability and not a sufficient one. The reason is that for $v$-representability of a density $\rho$ we additionally require that the variation over $\Gamma\mapsto\rho$ in Eq. (15b) has a minimum too, otherwise there is no Gibbs state behind the density. In this section, we show that the subdifferential of $F^{\beta}_{\mathrm{DM}}(\rho)$ at any density $\rho\in\mathscr{X}_{>0}$ is non-empty. In the next section we then show that these densities indeed come from Gibbs states. The two result together give then $v$-representability for all densities in $\mathscr{X}_{>0}$.
To show the existence of a subgradient for a density $\rho\in\mathscr{X}_{>0}$ we proceed as in Sutter et al. [7]. The only result that significantly changes compared to this reference is the local boundedness of the universal functional $F_{\mathrm{DM}}^{\beta}$.
Lemma 6.
There exist three constants $C_{1},C_{2},C_{3}>0$ such that for all $\rho\in\mathscr{I}$ it holds
$C_{1}\leq F_{\mathrm{DM}}^{\beta}(\rho)\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}.$ (18)
Proof.
For the lower bound of $F_{\mathrm{DM}}^{\beta}(\rho)$ we have
$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0)$ (19)
which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta S(\Gamma)$ is negative. Hence we get the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\Gamma\to\rho}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}$ (20)
where we used a previous result [7, Lem. 10] for the second inequality. ∎
Theorem 7.
For any $\rho\in\mathscr{X}_{>0}$ the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
The proof is the same as in Sutter et al. [7, Th. 15].
V Existence of the Minimum
As discussed in the previous section, we do not only need a non-empty subdifferential of $F_{\mathrm{DM}}^{\beta}(\rho)$ to get $v$-representability but also that the minimum in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ (Eq. (15b)) exists. This minimizer must then be the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$. In the following we show this for all $\rho\in\mathscr{X}_{>0}$. To do so, we make use of the relative entropy which is defined as
$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}$ (21)
for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$. An important property of the relative entropy is lower semi-continuity.
Lemma 8 (Lami and Shirokov [24], Lemma 4).
The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology. (for non-negative trace class operators).
The next lemma relates the Helmholtz functional and the relative entropy. Note that originally we only defined the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)$ for (normalized) density matrices. However, in the same manner it can be defined for any non-negative trace class operator. Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and it is given by $\mathrm{e}^{-\beta H_{v}}/Z(v)$.
Lemma 9 (Giesbertz and Ruggenthaler [16]).
$S(\Gamma|\Gamma_{v})+\operatorname{Tr}\{\Gamma-\Gamma_{v}\}=\beta\Omega^{\beta}_{v}(\Gamma)-\beta\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}$.
Since the left hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (norm is decreasing under weak-* limits), we get that the Helmholtz functional together with a term proportional to $\operatorname{Tr}\{\cdot\}$ is lower semi-continuous with respect to the weak-* topology. In particular, we get the following lemma.
Lemma 10.
For $\xi>1$ the mapping $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous if $\Omega^{\beta}_{v}(\Gamma_{v})\leq(\beta(\xi-1))^{-1}$. For $\xi<1$ the function is weak-* lower semi-continuous if $\Omega^{\beta}_{v}(\Gamma_{v})\geq(\beta(\xi-1))^{-1}$.
Proof.
From Lemma 9 we get
$\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}_{v}(\Gamma_{v})\operatorname{Tr}\{\Gamma\}=\frac{1}{\beta}S(\Gamma|\Gamma_{v})-\frac{1}{\beta}\operatorname{Tr}\{\Gamma_{v}\}+\operatorname{Tr}\{\Gamma\}\left(\frac{1}{\beta}+(1-\xi)\Omega^{\beta}_{v}(\Gamma_{v})\right).$ (22)
By Lemma 8 we know that the first term $\beta^{-1}S(\Gamma|\Gamma_{v})$ is weak-* lower semi-continuous. The second term $\beta^{-1}\operatorname{Tr}\{\Gamma_{v}\}$ is just a constant. The last term is weak-* lower semi-continuous if $\beta^{-1}+(1-\xi)\Omega^{\beta}_{v}(\Gamma_{v})\geq 0$. Solving for $\Omega^{\beta}_{v}(\Gamma_{v})$ proves the lemma. ∎
Theorem 11.
For any $\rho\in\mathscr{X}_{>0}$ there is a (normalized) density matrix $\Gamma\mapsto\rho$ such that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log\Gamma)\}$. This shows the existence of a minimizer in Eq. (15b).
Proof.
For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exists by Theorem 7. By definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ there is a sequence of density matrices $\Gamma_{n}$ each yielding the density $\rho$ and such that $\operatorname{Tr}\{\Gamma_{n}(H_{0}+\beta^{-1}\log\Gamma_{n})\}\to F^{\beta}_{\mathrm{DM}}(\rho)$. Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [18, Th. IV.21] that there is a subsequence (in the following also denoted by $\Gamma_{n}$) and a trace class operator $\Gamma_{0}$ such that $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$. Moreover, we know that $\Gamma_{0}$ is non-negative and that $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$. To prove the theorem, we show that $\Gamma_{0}$ is a minimizer. That means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ and that it yields the density $\rho$. We start by showing $\operatorname{Tr}\{\Gamma_{0}\}=1$.
For simplicity, we call the minimum of the Helmholtz functional $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$. Because $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
$\Omega^{\beta}-\xi\Omega^{\beta}$ $=F^{\beta}_{\mathrm{DM}}(\rho)+\langle v,\rho\rangle-\xi\Omega^{\beta}=\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log\Gamma_{n})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{n}\}$ (23)
$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
where we have used the fact that the function $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous due to Lemma 10. For the condition of the lemma to be fulfilled we set $\xi=(\beta\Omega^{\beta})^{-1}+1$. This expression is well-defined, since we can choose a representative of $v]\in\mathscr{X}^{*}$ for which $\Omega^{\beta}_{v}(\Gamma_{v})$ is non-zero. So the condition of Lemma 10 is satisfied. Let $c$ be the constant such that $\operatorname{Tr}\{c\Gamma_{0}\}=1$. (Note that the constant $c$ only exist if $\Gamma_{0}\neq 0$. If $\Gamma_{0}=0$ then the right hand-side of Eq. (23) is 0 where the left hand-side is $\beta^{-1}$. Hence $\Gamma_{0}=0$ is not possible.) Since $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$ it holds $c\geq 1$. Then we can compute
$\Omega^{\beta}-\xi\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ $=\operatorname{Tr}\{c\Gamma_{0}(H_{v}+\beta^{-1}\log(c\Gamma_{0}))\}-\xi\Omega^{\beta}\operatorname{Tr}\{c\Gamma_{0}\}$ (24)
$=c\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}+\beta^{-1}\operatorname{Tr}\{c\Gamma_{0}\}\log c-c\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
$=c(\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\})+\beta^{-1}\log c$
$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c$
where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq. (23). This inequality can be transformed to
$c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}\geq 0.$ (25)
It is clear that we have equality if $c$ is equal to $1$. The function $f(c)=c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}$ defined on $1,\infty)$ has $f(1)=0$ and the derivative
$f^{\prime}(c)=\Omega^{\beta}-\xi\Omega^{\beta}+\frac{\beta^{-1}}{c}=-\beta^{-1}+\frac{\beta^{-1}}{c}\begin{cases}=0\quad&c=1\\ <0\quad&c>1,\end{cases}$ (26)
where we used $\xi=(\beta\Omega^{\beta})^{-1}+1$. This means that $f(c)$ is strictly decreasing on $1,\infty)$ with a unique maximum at $c=1$. Hence the inequality (25) can only be satisfied if $c=1$ and we conclude that $\operatorname{Tr}\{\Gamma_{0}\}=1$.
Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\to\rho$. For this we use the same idea as Lieb [6] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$ where $M_{f}$ denotes pointwise multiplication by $f$. This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\to\rho$. To do this, let $|\sum_{j}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of $\Gamma_{0}$. Since $\operatorname{Tr}\{\Gamma_{0}\}=1$ there is for any $\epsilon>0$ an $N_{\epsilon}\in\mathbb{N}$ such that
$\sum\limits_{j=1}^{N_{\epsilon}}\lambda_{j}>1-\epsilon.$ (27)
We define the projection
$|.P=\sum\limits_{j=1}^{N_{\epsilon}}\ket{\psi_{j}}\bra{\psi_{j}}$. (28)
The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ in trace class, there exist $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$. It follows that for $n>N_{P}$ we have $\operatorname{Tr}\{\Gamma_{n}P\}>\operatorname{Tr}\{\Gamma_{0}P\}-\epsilon>1-2\epsilon$ and thus $\operatorname{Tr}\{\Gamma_{n}(1-P)\}<2\epsilon$. Then we get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$. We can write
$\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}=\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})PM_{f}\}+\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)M_{f}\}.$ (29)
The first term goes to 0 since $PM_{f}$ is a compact operator. The second term is contained in $C\epsilon,C\epsilon]$ for $n$ large enough where the constant $C$ depends only on $f$. Since $\epsilon$ can be arbitrarily small we conclude that $\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})M_{f}\}\to 0$ and hence $\Gamma_{0}\to\rho$. By Eq. (23) we get $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ which concludes the proof. ∎
Remark 1.
Note that we never use $\rho>0$ and that we can also replace the condition $\rho\in\mathscr{X}_{>0}$ by the condition that the subdifferential of $F_{\mathrm{DM}}^{\beta}$ at $\rho$ is non-empty.
If we combine Theorem 7 and Theorem 11, we get that each density $\rho\in\mathscr{X}_{>0}$ is $v$-representable, which is the first part of Theorem 1. The full proof for Theorem 1 can be found in Section VIII.
VI Closing the Circle
So far we showed that each density in $\mathscr{X}_{>0}$ is $v$-representable. In this section we show that each density coming from a Gibbs state is contained in $\mathscr{X}_{>0}$. At elevated temperatures, all excited state of the Hamiltonian contribute to the Gibbs state (and hence also to the density) and since these state form an orthonormal basis, it is not possible that the density vanishes somewhere. We formulate this result in the following theorem that closes the circle from strictly positive densities in $\mathscr{X}_{>0}$ to representing potentials in $\mathscr{X}^{*}$ and back to a strictly positive density of the corresponding Gibbs state.
Theorem 12.
The density of any Gibbs state is everywhere non-zero.
Proof.
Let $v\in H^{-1}(\mathbb{T})$. The Gibbs state is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ which can be written as
$|,\Gamma_{v}=\frac{1}{Z(v)}\sum_{j}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$, (30)
where the $\lambda_{j}$ are the eigenvalues of $H_{v}$ with corresponding eigenstates $\psi_{j}$. We denote the density of $\Gamma_{v}$ by $\rho$ and the densities of $\psi_{j}$ by $\rho_{j}$. Let us assume that the density $\rho$ is equal to 0 at $x_{0}$. By the relation
$\rho(x)=\frac{1}{Z(v)}\sum_{j}\mathrm{e}^{-\beta\lambda_{j}}\rho_{j}(x)$ (31)
we conclude that $\rho_{j}(x_{0})=0$ for all $j$ since the weights are all non-zero. This implies that all $\psi_{j}$ vanish on $U=\{x_{0}\}\times[0,1]^{N-1}$. (The restriction of $\psi_{j}$ to $U$ is well-defined by the trace operator for Sobolev spaces [25, p. 315].) Consequently, each linear combination of the $\psi_{j}$ vanishes on $U$. Yet, this statement makes only sense if the linear combination is in $H^{1}$ too. Since the $\psi_{j}$ form a basis of $L^{2}(\mathbb{T}^{N})$ we also need to be able to express each function in $H^{1}(\mathbb{T}^{N})\subset L^{2}(\mathbb{T}^{N})$ as their linear combination. Now this yields a contradiction since clearly not every function in $H^{1}(\mathbb{T}^{N})$ vanishes on U. We note that while the restriction to $U$ is well-defined for $H^{1}(\mathbb{T}^{N})$ functions, this is not the case for functions in $L^{2}(\mathbb{T}^{N})$. ∎
VII Gâteaux differentiability
The last part of the main result (Theorem 1) is about Gâteaux differentiability of $F_{\mathrm{DM}}^{\beta}$ at densities $\rho\in\mathscr{X}_{>0}$. As mentioned in Section IV, Gâteaux differentiability can be achieved by continuity and a single-element subdifferential. Thus, the goal of this section is to show that each density $\rho\in\mathscr{X}_{>0}$ has only one subgradient. We do that by using standard results from convex analysis and by exploiting the following lemma.
Lemma 13.
The Helmholtz free energy $\Omega^{\beta}(v)$ is a strictly concave function in $v\in H^{-1}(\mathbb{T})$.
This result is obtained by exploiting the uniqueness of the Gibbs state. The proof proceeds analogously to [16, Th. 10]. Additionally, we need the following definition and lemmata.
Definition 14 (Barbu and Precupanu [23], p. 71).
Let $f$ be a convex function on $\mathscr{X}$. The closure of $f$, denoted by $\mathrm{cl}f$, is the lower semi-continuous hull of $f$, i.e.  $\mathrm{cl}f(\rho)=\liminf_{\rho^{\prime}\to\rho}f(\rho^{\prime})$ for all $\rho\in\mathscr{X}$ if $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})>$-$\infty$ for all $\tilde{\rho}$ and it is the constant function -$\infty$ if there is a $\tilde{\rho}\in\mathscr{X}$ such that $\liminf_{\rho^{\prime}\to\tilde{\rho}}f(\rho^{\prime})=$-$\infty$.
Since the universal functional $F_{\mathrm{DM}}^{\beta}$ is uniformly bounded from below by Lemma 6, the closure of it is given by $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Another characterization of the closure of a convex function is by its biconjugate as the following lemma shows.
Lemma 15 (Barbu and Precupanu [23], Cor. 2.23).
The biconjugate of a convex function $f$ coincides with its closure. That is, $f^{**}=\mathrm{cl}f$.
Lemma 16 (Barbu and Precupanu [23], Prop. 2.33).
Let $f:\mathscr{X}\to\mathbb{R}\cup\{\infty\}$ be a proper, convex and lower semi-continuous function and $f^{*}$ its conjugate. Then the following two statements are equivalent,
  1. (i)
    $v\in\partial f(\rho)$
  2. (ii)
    $\rho\in\partial f^{*}(v)$.
Lemma 17 (Barbu and Precupanu [23], Prop. 2.40).
If a convex function $f$ on $\mathscr{X}$ is Gâteaux differentiable at $\rho$, then $\partial f(\rho)$ consists of a single element $v=\mathrm{grad}f(\rho)$. Conversely, if $f$ is continuous at $\rho$ and if $\partial f(\rho)$ contains only a single element, then $f$ is Gâteaux differentiable at $\rho$ and $\mathrm{grad}f(\rho)=\partial f(\rho)$.
We have now all the needed ingredients to show Gâteaux differentiability of the universal functional.
Theorem 18.
The universal functional $F_{\mathrm{DM}}^{\beta}$ is Gâteaux differentiable at any density $\rho\in\mathscr{X}_{>0}$.
Proof.
According to Theorem 7, the functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$. By Lemma 17 it is then enough to show that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued for all densities $\rho\in\mathscr{X}_{>0}$. The conjugate function of $F_{\mathrm{DM}}^{\beta}$ is given by $\Omega^{\beta}(-\cdot)$. If we conjugate this function we end up with the biconjugate of $F_{\mathrm{DM}}^{\beta}$ which is defined by
$F_{\mathrm{DM}}^{\beta**}(\rho)=\sup\big{\{}\Omega^{\beta}(v)-\langle\rho,v\rangle\mid v\in H^{-1}(\mathbb{T})\big{\}}.$ (32)
The biconjugate $F_{\mathrm{DM}}^{\beta**}$ is a convex, lower semi-continuous function and it coincides with the convex envelope of $F_{\mathrm{DM}}^{\beta}$. Since $F_{\mathrm{DM}}^{\beta}$ is convex, Lemma 15 states that the closure of $F_{\mathrm{DM}}^{\beta}$ equals to its biconjugate $F_{\mathrm{DM}}^{\beta**}$. By Definition 14 and its proceeding discussion $\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})$. Theorem 7 shows that the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous on $\mathscr{X}_{>0}$ and hence we get that for each $\rho>0$
$F_{\mathrm{DM}}^{\beta**}(\rho)=\mathrm{cl}F_{\mathrm{DM}}^{\beta}(\rho)=\liminf\limits_{\rho^{\prime}\to\rho}F_{\mathrm{DM}}^{\beta}(\rho^{\prime})=F_{\mathrm{DM}}^{\beta}(\rho).$ (33)
This means that $F_{\mathrm{DM}}^{\beta}$ and its biconjugate $F_{\mathrm{DM}}^{\beta**}$ coincide on the open set $\mathscr{X}_{>0}$. In particular this means that the subdifferentials are equal for densities $\rho\in\mathscr{X}_{>0}$, i.e. $\partial F_{\mathrm{DM}}^{\beta}(\rho)=\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. So instead of showing that $\partial F_{\mathrm{DM}}^{\beta}(\rho)$ is single-valued, we show it for $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$. Note that the conjugate function of $F_{\mathrm{DM}}^{\beta**}$ is $\Omega^{\beta}(-\cdot)$ since the triple conjugate of $F_{\mathrm{DM}}^{\beta}$ is equal to the conjugate. The function $F_{\mathrm{DM}}^{\beta**}$ fulfills all assumptions of Lemma 16 and hence we get
$v\in\partial F_{\mathrm{DM}}^{\beta**}(\rho)\Longleftrightarrow\rho\in\partial\left(-\Omega^{\beta}\right)(v).$ (34)
By Lemma 13 we have that $\Omega^{\beta}$ is strictly convex. In particular this means that a density $\rho$ can only be contained in one subdifferential. Otherwise, if $\rho$ is contained in the subdifferential at $v_{1}$ and $v_{2}$, then the function $\Omega^{\beta}$ would not be strictly convex on the line with boundary points $v_{1}$, $v_{2}$. By the relation (34) we get that $\partial F_{\mathrm{DM}}^{\beta**}(\rho)$ is single valued. This concludes the proof. ∎
VIII Summary and outlook
We can now collect the results from the previous sections to finally give a proof of our main result.
Proof of Theorem 1.
Let $\rho\in\mathscr{X}_{>0}$, then Theorem 7 shows that $\partial F^{\beta}_{\mathrm{DM}}(\rho)$ is non-empty. This implies that $\rho$ is a minimizer for the variational principle in Eq. (15a), where the negative potential is in the subdifferential, i.e., $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$. The minimizer of Eq. (15b), guaranteed to exist by Theorem 11, is then also a minimizer in the Helmholtz functional (Eq. (12)). This minimizer is unique by Proposition 3 and is equal to the Gibbs state $\Gamma_{v}$. We conclude, that $\rho$ is the density of $\Gamma_{v}$ and hence $v$-representable. On the other hand, every density of a Gibbs state $\Gamma_{v}$ with $v\in H^{-1}(\mathbb{T})$ is contained in $\mathscr{I}$ according to Proposition 4 and is strictly positive by Theorem 12. In particular, every $v$-representable density is in $\mathscr{X}_{>0}$ and $F^{\beta}_{\mathrm{DM}}$ is not differentiable at densities in $\mathscr{I}\setminus\mathscr{X}_{>0}$. On the other hand, Theorem 18 shows that the universal functional $F^{\beta}_{\mathrm{DM}}$ is Gâteaux differentiable at any $\rho\in\mathscr{X}_{>0}$. ∎
In this work we provided an explicit characterization of the set of uniquely $v$-representable densities on the one-dimensional torus at elevated temperatures and established (Gâteaux) differentiability of the universal functional. In the same manner the theory can be developed for any other one-dimensional bounded domain. As an extensive parameter, the volume needs to be finite, and thus a periodic setting is the natural choice for elevated temperatures. The restriction to one dimension remains as the main limitation of the current formulation. Another interesting observation is that we do need distributional potentials. In DFT often the potential set $L^{3/2}+L^{\infty}$ is considered [6]. However, for zero temperature and one particle it was already observed by Englisch and Englisch [26] and Chayes et al. [12] that relatively nice densities need distributional potentials. Sutter et al. [7] showed that distributional potentials are sufficient for $v$-representability of the set $\mathscr{X}_{>0}$. For elevated temperature we further show that distributional potential are also necessary. This leads to the clear conjecture that also for higher dimensional domains some sort of distributional potentials are needed. However, the continuous Sobolev embedding $H^{1}\hookrightarrow L^{\infty}$ is crucial and it holds only for one-dimensional domains. For higher dimension we either need to require higher regularity or higher integrability conditions. This means that for the Sobolev space $W^{m,p}$, we need to have $m$ and $p$ large enough. The problem is then that not all Gibbs state densities might be contained in $W^{m,p}$ and that a Hamiltonian with a potential in $W^{-m,q}=(W^{m,p})^{\ast}$ might not be self-adjoint. Another open question is how to extend the results to the grand canonical ensemble of variable particle number. In this case, the particle-number operator is not bounded anymore and this is currently needed in the proof of Theorem 11 where we show that the weak-* limit yields the same density as the sequence. Additionally, we do not know if the square root of the Gibbs state density is in $H^{1}$. However, we think that the rest of the theory does not fundamentally change and that we get the same conclusion as for the canonical ensemble.
Acknowledgements.
MP acknowledges support from the German Research Foundation under Grant SCHI 1476/1-1 and from ERC-2021-STG grant agreement No. 101041487 REGAL. The authors SMS and KJHG thank The Netherlands Organization for Scientific Research, NWO, for its financial support under Grant No. OCENW.KLEIN.434 and Vici Grant No. 724.017.001. RvL acknowledges the Academy of Finland grant under project number 356906. MR acknowledges the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994.
Appendix A Proof of Proposition 3
We still need to show that Proposition 3 holds. The result that the unique minimizer of the Helmholtz functional is given by the Gibbs state $\Gamma_{v}=\mathrm{e}^{\beta H_{v}}/Z(v)$ and the relation $\Omega^{\beta}_{v}(\Gamma_{v})=-\beta\log Z(v)$ is shown in standard literature. Yet, we give here a different approach than Mermin [11] and a more elaborated version than our previous works [16; 15].
Proof of Proposition 3.
To determine the minimizer of the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)=\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$ we make it stationary with respect to small perturbations $\delta\Gamma$ that obey $\operatorname{Tr}\{\delta\Gamma\}=0$. We compute
$\Omega^{\beta}_{v}(\Gamma+\epsilon\delta\Gamma)-\Omega^{\beta}_{v}(\Gamma)=\epsilon\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{(\Gamma+\epsilon\delta\Gamma)\log(\Gamma+\epsilon\delta\Gamma)\}-\beta^{-1}\operatorname{Tr}\{\Gamma\log\Gamma\}.$ (35)
We use the Taylor expansion of $x\log(x)$ to get $(\Gamma+\epsilon\delta\Gamma)\log(\Gamma+\epsilon\delta\Gamma)=\Gamma\log\Gamma+\log(\Gamma)\epsilon\delta\Gamma+\epsilon\delta\Gamma+\mathcal{O}(\epsilon^{2})$. Dividing by $\epsilon$ turns Eq. (35) into
$\frac{\Omega^{\beta}_{v}(\Gamma+\epsilon\delta\Gamma)-\Omega^{\beta}_{v}(\Gamma)}{\epsilon}$ $=\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\}+\mathcal{O}(\epsilon)$ (36)
$=\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma\}+\mathcal{O}(\epsilon).$
The quotient in Eq. (36) must go to zero for any minimizer $\Gamma_{v}$ as $\epsilon\to 0$. Hence we have $\operatorname{Tr}\{\delta\Gamma H_{v}\}+\beta^{-1}\operatorname{Tr}\{\delta\Gamma\log\Gamma_{v}\}=\operatorname{Tr}\{\delta\Gamma(H_{v}+\beta^{-1}\log\Gamma_{v})\}=0$ for all $\delta\Gamma$ which requires $H_{v}+\beta^{-1}\log\Gamma_{v}=\mathrm{const}$. This then leads to the unique minimizer $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)$ is determined from the normalization constraint.
To get the expression for the Helmholtz free energy, let us first assume that $Z(v)$ is finite, i.e., that the Gibbs state exists. Using the explicit expression of the Gibbs state $\Gamma_{v}$ and $\log\Gamma_{v}=-\beta H_{v}-\log Z(v)$ gives the expression
$\Omega^{\beta}(v)=\Omega^{\beta}_{v}(\Gamma_{v})=\operatorname{Tr}\{\Gamma_{v}(H_{v}+\beta^{-1}\log\Gamma_{v})\}=\operatorname{Tr}\{\Gamma_{v}(H_{v}-H_{v}-\beta^{-1}\log Z(v))\}=-\beta^{-1}\log Z(v).$ (37)
For the last remaining part of the proposition we need to show that $Z(v)$ is indeed finite for any potential $v\in H^{-1}(\mathbb{T})$. First, we show that the partition function is finite for a non-interacting system with zero potential, i.e., $H_{v}=T$. We denote this partition function by $Z_{0}(0)$. The eigenstates of the kinetic-energy operator $\frac{1}{2}\Delta$ on $\mathbb{T}$ are Slater determinants built from the functions $\mathrm{e}^{\mathrm{i}2\pi p\cdot x}$ with $p\in\mathbb{Z}$ which have eigenvalues $2\pi^{2}p^{2}$. We estimate
$Z_{0}(0)=\operatorname{Tr}\{\mathrm{e}^{-\beta T}\}\leq\sum\limits_{p_{1},\ldots,p_{N}\in\mathbb{Z}}\mathrm{e}^{-\beta 2\pi^{2}(p_{1}^{2}+\ldots+p_{N}^{2})}=\Big{[}\sum\limits_{p\in\mathbb{Z}}\mathrm{e}^{-\beta 2\pi^{2}p^{2}}\Big{]}^{N}<\infty.$ (38)
To show that $Z(v)$ is finite for a general interaction and a general potential we first show that $\Omega^{\beta}(v)$ is finite. For that recall that we assume that the interaction satisfies the KLMN condition (3). Moreover, Sutter et al. [7, Lem. 16] showed that any potential $v\in H^{-1}$ fulfills also the KLMN condition $|V(\Psi,\Psi)|\leq aT(\Psi)+b$ where $a>0$ can be made arbitrary small. Here $V(\cdot,\cdot)$ is the quadratic form produced by $v\in H^{-1}$. Combing these two KLMN conditions guarantees the existence of $0<a<1$ and $b$ such that $\langle\Psi,H_{v}\Psi\rangle\geq(1-a)T(\Psi)-b$ for all $\Psi\in\mathscr{W}$. This condition translates to density matrices and we get
$\Omega^{\beta}(v)=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log(\Gamma))\}\geq(1-a)\inf_{\Gamma}\operatorname{Tr}\{\Gamma(T+\big{(}\beta(1-a)\big{)}^{-1}\log(\Gamma))\}-b=-\beta^{-1}\log\big{(}Z_{0}(0)\big{)}-b>-\infty.$ (39)
Let $|H_{v}=\sum_{j=1}^{\infty}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of the Hamiltonian $H_{v}$ which exist due to the compact embedding of $Q(H_{v})=\big{(}H^{1}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$ into $\big{(}L^{2}(\mathbb{T})\otimes\mathbb{C}^{2}\big{)}^{\wedge N}$. We define a sequence of density matrices by
[…]
and we compute
$\Omega^{\beta}(v)<\operatorname{Tr}\{\Gamma_{n}(H_{v}+\beta^{-1}\log(\Gamma_{n}))=\frac{1}{Z_{n}[v]}\sum\limits_{j=1}^{n}\mathrm{e}^{-\beta\lambda_{j}}(-\beta^{-1}\log Z_{n}[v])=-\beta^{-1}\log Z_{n}[v].$ (41)
The left-hand side is bounded from below by Eq. (39) and we get that $Z_{n}[v]$ is uniformly bounded from above. This implies $Z(v)<\infty$. ∎
Appendix B Proof of Proposition 4
In this section we show that the density of a Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$. To do so, note that as explained in the proof of Proposition 3 we have for each potential $v$ two constants $0\leq a<1$ and $b\geq 0$ such that
$(1-a)T-b\leq H_{v}\leq(1+a)T+b$ (42)
where $T=-1/2\Delta$ is the kinetic energy operator. Moreover, the spectrum of the Hamiltonian $H_{v}$ (and also of $T$) contains only discrete values due to the compact embedding $Q(H_{v})=H^{1}(\mathbb{T}^{N})\hookrightarrow L^{2}(\mathbb{T}^{N})$. Then we can express $H_{v}$ and $T$ in spectral decomposition
[…]
where we additionally require $\lambda_{1}\leq\lambda_{2}\leq\ldots$ and $\mu_{1}\leq\mu_{2}\leq\ldots$. Then, we get the following lemma.
Lemma 19.
The $n$’th eigenvalue of $H_{v}$ has the lower and upper bound, $(1-a)\mu_{n}-b\leq\lambda_{n}\leq(1+a)\mu_{n}+b$.
Proof.
Let $\phi_{j}$ be the states introduced in Eq. (43). By the min-max-principle for self-adjoint operators and by Eq. (42) we get
[…]
For the lower bound we proceed in the same way but we use the max-min-principle instead to get
[…]
We use the lower and upper bound of the eigenvalues to proof that the kinetic energy expectation is finite.
Lemma 20.
For all, $v\in H^{-1}$ we have $\mathrm{Tr}\{T\Gamma_{v}\}<\infty$.
Proof.
The Gibbs state is given by $|\Gamma_{v}=\frac{1}{Z(v)}\sum_{j=1}^{\infty}\mathrm{e}^{-\beta\lambda_{j}}\ket{\psi_{j}}\bra{\psi_{j}}$. We first show that the energy expectation is finite by using Lemma 19. That is,
-$\infty$ $<\mathrm{Tr}\{H_{v}\Gamma_{v}\}=\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\lambda_{j}\mathrm{e}^{-\beta\lambda_{j}}\leq\frac{1}{Z(v)}\sum\limits_{j=1}^{\infty}\big{[}(1+a)\mu_{j}+b\big{]}\mathrm{e}^{-\beta[(1-a)\mu_{j}-b]}$
$=\frac{(1+a)}{Z(v)}\mathrm{e}^{\beta b}\sum\limits_{j=1}^{\infty}\mu_{j}\mathrm{e}^{-\beta(1-a)\mu_{j}}+\frac{\mathrm{e}^{\beta b}b}{Z(v)}\sum\limits_{j=1}^{\infty}\mathrm{e}^{-\beta(1-a)\mu_{j}}.$
A similar computation as in Eq. (38) shows that the two sums are finite and we get $\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty$. Using Eq. (42) and positivity of the kinetic energy operator gives
$-b\leq(1-a)\mathrm{Tr}\{T\Gamma_{v}\}-b\leq\mathrm{Tr}\{H_{v}\Gamma_{v}\}<\infty,$
which concludes the proof. ∎
We can now prove Proposition 4.
Proof of Proposition 4.
For simplicity we introduce $\alpha_{j}=\mathrm{e}^{-\beta\lambda_{j}}/Z(v)$ such that the Gibbs state is $|.\Gamma_{v}=\sum_{j=1}^{\infty}\alpha_{j}\ket{\psi_{j}}\bra{\psi_{j}}$. Then, its density is given by $\rho=\sum_{j=1}^{\infty}\alpha_{j}\rho_{j}$ with $\psi_{j}\mapsto\rho_{j}$. By Theorem 2 all the pure state densities $\rho_{j}$ are contained in $\mathscr{I}$. We define a sequence $\rho^{n}=\sum_{j=1}^{n}\alpha_{j}\rho_{j}$. Clearly, $\rho^{n}\to\rho$ in $L^{1}(\mathbb{T})$ and $\|\sqrt{\rho^{n}}\rVert_{2}^{2}\leq N$. Since $\int(\nabla\sqrt{\rho})^{2}\mathrm{d}x$ is convex on $\mathscr{I}$ [6, remark after Eq. (1.10)], we find by Jensen’s inequality and by Theorem 2
$\int(\nabla\sqrt{\rho^{n}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}\int(\nabla\sqrt{\rho_{j}})^{2}\mathrm{d}x\leq\sum\limits_{j=1}^{n}\alpha_{j}2T(\psi_{j})\leq 2\mathrm{Tr}\{T\Gamma_{v}\}<\infty,$
where we get a uniform bound by Lemma 20. In particular, $\sqrt{\rho^{n}}$ is a bounded sequence in $H^{1}(\mathbb{T})$ and by the Banach–Alaoglu theorem, there is a subsequence (also denoted by $\sqrt{\rho^{n}}$) and $f\in H^{1}(\mathbb{T})$ such that $\sqrt{\rho^{n}}\rightharpoonup f$ weakly in $H^{1}$. But since $H^{1}(\mathbb{T})$ is compactly embedded in $L^{2}(\mathbb{T})$ there is also a subsequence (again denoted by $\sqrt{\rho^{n}}$) such that $\sqrt{\rho^{n}}\to f$ strongly in $L^{2}(\mathbb{T})$. Then, we have
$\|\rho^{n}-f^{2}\rVert_{1}=\int(\sqrt{\rho^{n}}-f)(\sqrt{\rho^{n}}+f)\mathrm{d}x\leq\|\sqrt{\rho^{n}}-f\rVert_{2}\|\sqrt{\rho^{n}}+f\rVert_{2}\to 0$
since $\|\sqrt{\rho^{n}}+f\rVert_{2}$ is bounded. This means that $\rho^{n}\to f^{2}$ strongly in $L^{1}$ and we conclude $\sqrt{\rho}=f\in H^{1}(\mathbb{T})$. ∎